OpenStudy (anonymous):

What's the value of tan (-2pi)/3?

OpenStudy (kendricklamar2014):

\[\frac{ \tan(−2)(\pi) }{ 3 } = \frac{ \tan(−6.283185...) }{ 3 }\]

OpenStudy (kendricklamar2014):

\[\frac{ \tan(−6.283185...) }{ 3 } = \frac{ 0 }{ 3 }\]

OpenStudy (anonymous):

my options are: * 1/2 * sqrt3 * -1/sqrt3 * sqrt 3/2

OpenStudy (kendricklamar2014):

\[\frac{ 0 }{ 3 } = 0\] \[Answer: 0\]

OpenStudy (anonymous):

what went wrong?

OpenStudy (kendricklamar2014):

Nothing, the answer is Choice D

OpenStudy (anonymous):

medal given :)

OpenStudy (kendricklamar2014):

:)